OpenStudy (anonymous):

alculate the de Broglie wavelength of a 143-g baseball traveling at 92mph .

OpenStudy (anonymous):

lambda = h /(root(2*Ke*m) where: lambda = wavelenght h = planck constant Ke = kinetic energy ((1/2) mv2) m = mass